# Leveraging Self-Supervised Learning for Scene Recognition in Child Sexual Abuse Imagery

### v1

@@ Abstract
Crime in the 21st century is split into a virtual and real world. However, the former has become a global menace to people’s well-being and security in the latter. The challenges it presents must be faced with unified global cooperation, and we must rely more than ever on automated yet trustworthy tools to combat the ever-growing nature of online offenses. Over 10 million child sexual abuse reports are submitted to the US National Center for Missing & Exploited Children every year, and over 80% originated from online sources. Therefore, investigation centers and clearinghouses cannot manually process and correctly investigate all imagery. In light of that, reliable automated tools that can securely and efficiently deal with this data are paramount. In this sense, the scene recognition task looks for contextual cues in the environment, being able to group and classify child sexual abuse data without requiring to be trained on sensitive material. The scarcity and limitations of working with child sexual abuse images lead to self-supervised learning, a machine-learning methodology that leverages unlabeled data to produce powerful representations that can be more easily transferred to target tasks. This work shows that self-supervised deep learning models pre-trained on scene-centric data can reach 71.6% balanced accuracy on our indoor scene classification task and, on average, 2.2 percentage points better performance than a fully supervised version. We cooperate with Brazilian Federal Police experts to evaluate our indoor classification model on actual child abuse material. The results demonstrate a notable discrepancy between the features observed in widely used scene datasets and those depicted on sensitive materials.
@@ I Introduction
With the popularization of the internet, the modern world has been deemed to exist without physical borders, undeniably changing how we obtain information, communicate, work, consume, and socialize, but also providing an environment for misuse and abuse. New vicious forms of crime include spreading malicious software, global networks of abuse imagery, drug trafficking, illegal organ supply chains, and even artificial intelligence (AI) enabled blackmail. As crime develops and changes in the age of connected societies, so must policing, for “the world is becoming a single jurisdiction” [1] and efforts to fight crime in the 21st century must be made internationally. The spread of Child Sexual Abuse Materials (CSAM) is one such crime that has reached alarming proportions in the digital age. In the United States alone, the National Center for Missing and Exploited Children (NCMEC) reports a record 32 million reports of suspected CSAM in 2022 in the US [^1] https://www.missingkids.org/cybertiplinedata. The availability and sharing of such harmful content online not only exacerbate the trauma inflicted upon the victims but also create a significant burden on law enforcement agents who have to inspect thousands of files, leading to emotional strain manually [2].
In this paper, we look into the challenges of automating Child Sexual Abuse Imagery (CSAI) detection to present a new under-explored dimension of the problem – Indoor Scene Recognition – and thoroughly study how state-of-the-art foundation models [3] perform under this task, both on publicly available datasets and CSAI directly. With partner law enforcement agents, we design and present new class splits on public scene datasets, allowing for the evaluation of different methods and hyperparameter tuning without accessing CSAI beyond the final classification test carried out by the agents.
@@
Solutions that use visual clues to detect novel CSAI have not been absent but have focused on two main dimensions of what is expected of these materials from the outside: age estimation and nudity detection [7, 8, 9, 10, 11]. The recent study from Laranjeira da Silva et al. [12] has found, however, that these are not the only aspects that can be explored without direct data access; they highlight how, in the interviews with law enforcement agents performed by Kloess et al. [13, 14] and Yiallourou et al. [15], the illumination, gestures, facial expressions, and scene context are just as valuable to the classification of the material as legal or not.
@@
With this expanded view of the CSAI automation problem and the goal of adding to the tools available to law enforcement, we introduce the first Indoor Scene Classification method designed for use on CSAI; in developing this task, we yield both a tool for the triage of suspect materials but also a better understanding of how well the knowledge of methods trained public scene datasets transfer (or not) to CSAI and how this can be improved. The choice of specifically Indoor Scenes comes from talks with partner agents (the crime is seldom recorded in public spaces) and the findings of other studies. For instance, detecting multiple images of a child’s room potentially indicates the presence of children’s images in a database, even though they are not naked or undergoing any sexually explicit activity [2]. Indoor Scene Classification is also attractive for its feasibility with reliable, abuse-free material available in large quantities for training.
@@
Given that this is a novel task, we present a thorough investigation of the performance of foundation models, specifically models trained through self-supervision [16], when fine-tuned on Indoor Scene Classification as a downstream task. The task of self-supervision – learning without labels – is driven by the desire for more general representations that can be easily fine-tuned for downstream tasks unrelated to the original dataset, a goal well aligned with our CSAI classification needs. We do find, however, that because most self-supervised models are trained on object-centric datasets (namely ImageNet [17]), knowledge is not as easily transferable to a scene-centric task and that there are benefits to be had from doing self-supervision with scenes in addition to objects. Our set of contributions is then multi-faceted, from the introduction of indoor scene recognition to its application on CSAI data:
@@
    We introduce the task of Indoor Scene Classification, with new data splits on classic datasets tailored to yield models later useful for CSAI;
@@
    Our thorough evaluation of self-supervised models shows both their usefulness for Indoor Scene Recognition but also how further self-supervision on actual indoor scenes improves their downstream performance;
@@
The remaining sections are organized as follows. In Sections II-A and II-B, we review self-supervised and contrastive learning frameworks under deep learning. In Sections II-C and II-D, we bring the most recent works on scene recognition and CSAI investigation. In Section III, we describe the datasets, general pipelines used for pretraining, and fine-tuning models used in this work. Then, in Section V, we present the results obtained for our target task and real CSAI data. Finally, in Section VI, we point out the contributions, challenges, and possible routes for future work.
@@ II-A Self-supervised Learning
Self-supervised learning (SSL) is a machine-learning approach for building models using unlabeled data. Using unannotated data reduces data annotation costs while finding a purpose to the massive amounts of unlabeled images and texts published daily worldwide [18]. SSL follows a pattern similar to supervised transfer learning with a pre-training step called pretext task, designed to extract features from unlabeled data. In computer vision, common pretext tasks are grayscale colorization [19], jigsaw puzzle, cutout reconstruction, image inpainting, resolution upscaling, foreground object segmentation, clustering, temporal order verification, visual-audio correspondence verification, and contrastive representation comparison.
Afterward, the resulting model is fine-tuned on specific labeled downstream tasks without large amounts of data as a requirement. There are no limitations to the domain of the downstream task, but it has been empirically shown that ideally the pretext and downstream tasks should comprise images with similar properties, such as textures, shapes, and color gamut [20]. It has been demonstrated that self-supervised models outperform their supervised counterparts for ResNet-50 on image classification, detection, and segmentation in multiple benchmarks [20]. Therefore, SSL shows real advantages over supervised learning, given its higher robustness and generally better performance, especially for CSAI data, which has less than a hundred thousand images in the dataset [21].
@@ II-B Contrastive Learning
Contrastive learning, on the other hand, was thus developed as a new paradigm for self-supervised pre-training. It is closely related to metric learning, an unsupervised learning sub-field that aims to construct a specialized distance metric from weakly supervised data [22].
In the end-to-end contrastive learning framework, two encoders are trained simultaneously with contrastive loss. This method focuses on creating a model that, ideally, compares two samples numerically (e.g., two images), therefore grouping similar images while pushing different ones apart. Within an unlabeled dataset, there is no prior distance measurement. Thus, any augmented version is labeled as “similar” (or positive) for an anchor image sample, while all other samples are labeled as “different” (or negative). The first methods to propose this kind of instance discrimination were IntDisc [23], CMC [24], and Deep InfoMax [25]. However, it was notoriously successful with [26], which best demonstrated the importance of hard positive samples by introducing a set of 10 data augmentations from which transformations are sampled.
Despite its state-of-the-art performance on several downstream tasks, the end-to-end training is computationally expensive and does not always lead to the best downstream performance [20]. Thus, many other approaches were developed, such as MoCo [27, 28], SwAV [29], and Barlow Twins [30] to deal with batch size sensitivity, BYOL [31] and SimSiam [32] to overcome the need for negative sampling or SupCon [33] to use labels in SSL.
II-C Scene Recognition
Humans analyze their environment and important events efficiently, but machines have long stayed caught up in this topic. Scenes are understood as semantically coherent views of real-world environments containing objects, textures, and a spatially separated background [34, 35]. Scene recognition is a well-studied field with a particular interest in surveillance, autonomous driving, and robotics navigation, but it is still lacking in performance [36].
Scene classification is closely related to object and texture classification but with multiple objects, textures, backgrounds, and intricate relations between them. Thus, one could imagine it as multiple local classifications happening simultaneously to provide a context and a set of items or actions [37]. However, such a process would be practically impossible to program directly, yet current solutions [18] use deep learning and scene representation.
@@ II-C Scene Recognition
Additionally, Qiu et al. [41] used graph nets and image inpainting to show that a scene constitutes minor and essential objects, and only the essential ones matter for the final scene classification while minor elements can be removed without impact. However, the number of minor objects is usually much higher and can be regarded as noise.
@@
After selecting a pre-trained model, most works develop specialized fine-tuning routines to extract the most scene-centric features in the target dataset [37]. GAP-CNN replaces fully connected layers for global average pooling (GAP) to focus on class-specific regions while using fewer parameters [42]; Hierarchical LSTM uses four LSTM modules in an attempt to capture spatial dependencies [43]; DL-CNN proposes dictionary learning (DL) layers to obtain sparse representations and reduce the total number of parameters; finally, the Spatially Unstructured layer was designed to help cope with layout deformations, and scale changes in scene recognition [44]. Fisher Vector (FV-CNN) [45] and Mixture of Factor Analyzers Fisher Vector (MFAFVNet) [46] use Improved FV while Multi-scale Orderless Pooling (MOP-CNN) [47] and Semantic Descriptor with Objectness (SDO) [48] employ a Vector of Locally Aggregated Descriptors (VLAD) to cluster local features. However, such an approach heavily increases the number of images to be processed.
@@
Moreover, many studies [49, 50, 51, 52] used object detectors [53, 54, 55, 56, 57] to help with identification of essential scene regions and forcing the model to pay attention to particular objects during classification. Typically, this dependency introduces two-stage training [58], meaning that learning has to be split between the object detector and classifier.
@@
Several scene classification methods have used multiple layers’ outputs to improve scene representation in this context. The high CNN layers are too compact for the dense level features required for scene recognition, keeping only large objects, while the low CNN layers still keep the small ones [59]. Thus, harnessing multi-layer features reaches overall better representations, but careful feature fusion design is required [60]. Ultimately, this is the most structurally complicated approach, and the fused features generate high-dimensional arrays, making the model overfit easily the more layers are used [61].
@@
To the best of our knowledge, few works have used SSL as means to solve scene classification. In scene semantic segmentation, McCormac et al. [62] showed an improved model on NYUv2 and SUN RGB-D benchmarks using a model pre-trained on a 5M image 3D rendered dataset called SceneNet RGB-D. In order to tackle the PASCAL VOC benchmark (11,530 images containing 27,450 ROI annotated objects and 6,929 segmentations), Ren and Lee [63] proposed a multi-task learning approach that pre-trains an AlexNet model to predict edges, surface normals, and depth maps from SceneNet RGB-D images and a discriminator to distinguish the synthetic images from the Places365 images. Similarly, She and Xu [64] added a SimCLR contrastive loss as a new pretext task, setting as positive pairs all viewpoints of a single scene. The aforementioned methods reached comparable performance or even outperformed other SSL techniques using the same neural network architecture but underperformed compared to the state-of-the-art method for the tested benchmarks.
@@ II-D CSAI Classification
Easy access and impunity led to a frenzied growth in the consumption and distribution of CSAI over the past few years, with hundreds of millions of images circulating through the web today [2]. The automation of CSAI recognition can protect adults and children, more easily blocking that sort of content in social media and thus preventing trauma [4], but this is still a challenging field of research. Such datasets must stay accessible exclusively to law enforcement personnel, dramatically increasing the difficulties in comparing possible models or performing benchmark evaluations [12]. NCMEC shows that hash comparison tools are already the most significant source for reports today, but law enforcement personnel are the exclusive mechanism responsible for auditing CSAI when dealing with never-before-seen data, which is not only the source of bias but also psychological strain on the agents’ mental health [7].
@@
In light of these issues, researchers have attempted to design more scalable and reliable methodologies to tackle CSAI recognition. Image hash databases, web crawlers, and file metadata are most commonly used to detect CSAI sources and find criminals [65, 66, 67]. In a more versatile method, NuDetective [68] and iCOP [69] make use of handcrafted nudity image descriptors to build fast CSAI detectors, and Sae-Bae et al. [70] used textures and facial distances to distinguish adult from child nudity. However, these solutions are not robust against simple modifications, demanding constant updates to stay functional.
@@
With the recent developments in computer vision detectors, machine learning algorithms also started to be used for CSAI recognition. Ground-up training and transfer learning of CNNs were shown to outperform current forensic and commercial tools easily and can be made into portable tools for search and seizure procedures [8]. Moreover, Gangwar et al. [9] demonstrated that pornography and age-group recognition could be leveraged for CSAI with substantial gains in accuracy for binary classification, and Macedo et al. [7] used a single-model estimation of child presence, age, and gender to improve the performance. Rondeau [71] arguments apparent age and nudity detection could also be leveraged for CSAI, while Castrillón-Santana et al. [10] focused on non-adult people detection in images for CSAI age estimation, reaching over 90% accuracy on their proposed dataset.
@@
All these studies firmly outline that deep learning can contribute to child abuse inquiries. However, most published works focus on detecting children and nudity/pornography [72, 73, 74] and highlight the difficulties of finding large annotated CSAI databases to experiment with, even with the help of law enforcement agents. Several works focus exclusively on age estimation within CSAI, and children’s images are collected to build datasets that help accomplish such tasks. For instance, Anda et al. [75] proposed VisAGe and Castrillón-Santana et al. [10] AgeMega, databases focused on age estimation with thousands of underage and adult facial images, while [9, 10, 11] gather images from several age estimation databases. Beyond that, Al-Nabki et al. [76] and Tabone et al. [77] included reproductive organ detection to aid CSAI detection, while Yiallourou et al. [78] used facial expressions to distinguish sensitive from non-sensitive material, which is one of the main challenges in the field. Lee et al. [4] call attention to the fact that “the best results can be achieved if multiple methods are used in combination”, meaning that CSAI is a complex enough problem that age, nudity, and pornography detectors do not fulfill all the roles for possible feature extractors that can aid on this task.
@@
Notwithstanding these efforts, CSAI classification presents several challenges before accurate predictions can be made automatically. As Macilotti [79] explains, the decision of whether an image is explicit enough to be considered indecent is not well defined: the notion of “sexually explicit” may sound straightforward but becomes extremely subtle more often than not, given the context of the image or the culture of the viewers. Kloess et al. [80] argue that “nudity alone is not indicative of indecency” whereas posing fully clothed can be considered erotic under some circumstances, mainly if other images depicting the same child in more abusive material are found during the investigation.
@@
Moreover, Laranjeira da Silva et al. [12] designed an aggregated analysis pipeline to aid CSAI recognition research by showing a general pipeline for statistical analysis of features, such as scenes, skin tone, and nudity, making use of the region-based annotated child pornography dataset (RCPD) [7], without publicizing information on individual samples. RCPD is a private database used internally by the Brazilian Federal Police and contains 2,138 images among CSAI, pornography, and non-sensitive categories.
@@
In this context, most databases used for CSAI recognition could suffer from biased distributions, not representing what is found in the real world. Laranjeira da Silva et al. [12] show that CSA data shared online differs from reported cases of physical abuse in Brazil in terms of race: “most reported victims of child sexual abuse in Brazil are black and brown girls from 8 to 14 years old, but RCPD depicts mostly white children being abused and most of the material apprehended in Brazil has different tendencies than reports of sexual abuse with physical contact”.
@@
Overall, CSAI classification could take advantage of non-direct CSAI recognition, and Bursztein et al. [2] emphasize that the investigation should optimally consider scene information as means to cluster similar imagery and “report objects that are present, its environment, and potentially identify landmarks that will help locate the region where an abusive image originated”. Experts use scene information to gather context and understand what is happening in the image or video, especially in low-resolution scenarios and when people are covered or with their backs turned [80]. CSAI distribution is a massive problem worldwide, and, for the sake of the children under abuse, combating it demands new data-driven scalable solutions.

### v2

# Leveraging Self-Supervised Learning for Scene Classification in Child Sexual Abuse Imagery

@@ Abstract
Crime in the 21st century is split into a virtual and real world. However, the former has become a global menace to people’s well-being and security in the latter. The challenges it presents must be faced with unified global cooperation, and we must rely more than ever on automated yet trustworthy tools to combat the ever-growing nature of online offenses. Over 10 million child sexual abuse reports are submitted to the US National Center for Missing & Exploited Children every year, and over 80% originate from online sources. Therefore, investigation centers cannot manually process and correctly investigate all imagery. In light of that, reliable automated tools that can securely and efficiently deal with this data are paramount. In this sense, the scene classification task looks for contextual cues in the environment, being able to group and classify child sexual abuse data without requiring to be trained on sensitive material. The scarcity and limitations of working with child sexual abuse images lead to self-supervised learning, a machine-learning methodology that leverages unlabeled data to produce powerful representations that can be more easily transferred to downstream tasks. This work shows that self-supervised deep learning models pre-trained on scene-centric data can reach 71.6% balanced accuracy on our indoor scene classification task and, on average, 2.2 percentage points better performance than a fully supervised version. We cooperate with Brazilian Federal Police experts to evaluate our indoor classification model on actual child abuse material. The results demonstrate a notable discrepancy between the features observed in widely used scene datasets and those depicted on sensitive materials.
@@ I Introduction
With the popularization of the internet, the modern world has been deemed to exist without physical borders, undeniably changing how we obtain information, communicate, work, consume, and socialize, but also providing an environment for misuse and abuse. New vicious forms of crime include spreading malicious software, global networks of abuse imagery, drug trafficking, illegal organ supply chains, and even artificial intelligence (AI) enabled blackmail. As crime develops and changes in the age of connected societies, so must policing, for “the world is becoming a single jurisdiction” [1] and efforts to fight crime in the 21st century must be made internationally. The spread of Child Sexual Abuse Materials (CSAM) is one such crime that has reached alarming proportions in the digital age. In the United States alone, the National Center for Missing and Exploited Children (NCMEC) reports a record 32 million reports of suspected CSAM in 2022 in the US [^1] https://www.missingkids.org/cybertiplinedata. The availability and sharing of such harmful content online not only exacerbate the trauma inflicted upon the victims but also create a significant burden on law enforcement agents who have to inspect thousands of files, leading to emotional strain when done manually [2, 3].
In this paper, we look into the challenges of automating Child Sexual Abuse Imagery (CSAI) detection to present a new under-explored dimension of the problem — Indoor Scene Classification — and thoroughly study how state-of-the-art foundation models [4] perform under this task, both on publicly available datasets and CSAI directly. With partner law enforcement agents, we design and present new class splits on public scene datasets, allowing for evaluating different methods and hyperparameter tuning without accessing CSAI beyond the final classification test carried out by the agents.
@@
Solutions that use visual clues to detect novel CSAI have not been absent but have focused on two main dimensions of what is expected of these materials from the outside: age estimation and nudity detection [3, 8, 9, 10, 11, 12]. The recent study from [13] has found, however, that these are not the only aspects that can be explored without direct data access; they highlight how, in the interviews with law enforcement agents performed by [14, 15] and [16], the illumination, gestures, facial expressions, and scene context are just as valuable to the classification of the material as legal or not.
@@
With this expanded view of the CSAI automation problem and the goal of adding to the tools available to law enforcement, we introduce the first Indoor Scene Classification method designed for use on CSAI; in developing this task, we yield both a tool for the triage of suspect materials but also a better understanding of how well the knowledge of methods trained with public scene datasets transfer (or not) to CSAI and how this can be improved. The choice of specifically Indoor Scenes comes from talks with partner agents (the crime is seldom recorded in public spaces) and the findings of other studies. For instance, detecting multiple images of a child’s room potentially indicates the presence of children’s images in a database, even though they are not naked or undergoing any sexually explicit activity [2]. Indoor Scene Classification is also attractive for its feasibility with reliable, abuse-free material available in large quantities for training.
Given that this is a novel task, we present a thorough investigation of the performance of foundation models, specifically models trained through self-supervision [17], when fine-tuned on Indoor Scene Classification as a downstream task.
We aim to study the performance of self-supervised methods on our Indoor Scene Classification task and apply the best model on CSAI following the experimental pipeline at Fig. 1. Our goal is to use the Places8 dataset — created with law enforcement guidance — to approximate performance in CSAI. Because sensitive data is restricted in its access, we do not hope to use it for more than one evaluation with a single model.
We henceforth consider the problem through the lenses of three research questions, the first two about the data used to pre-train the model and the last about the method:
  1. Q1
    Does performing SSL on an object-centric dataset help or hurt downstream performance? Following the findings [18] that showed the combination of training on scene and object-centric datasets was beneficial to both tasks, we consider how SSL on ImageNet (an object-centric dataset) impacts the downstream performance on Places8. This is in addition to our comparison against a supervised baseline model pre-trained on ImageNet and fine-tuned on Places8.
  2. Q2
  3. Q3
With these questions and the experiments that will answer each in mind, our analysis can be summarized by the model variants yielded from the combination of the choices enlisted through the questions. When comparing these models, we use Bayesian estimation — replacing the common t-test — for ranking and contrasting. Bayesian Estimation Supersedes the T-test (BEST) [23] offers a nuanced and comprehensive view of experimental results compared to traditional t-tests. In our particular case, using 5-fold cross-validation breaks the independence assumption of the traditional t-test, leading to unreliable results. Also, we measure the advantage of using SSL over supervised techniques, which is accurately done using Bayesian methods.
@@
The task of self-supervision — learning without labels — is driven by the desire for more general representations that can be easily fine-tuned for downstream tasks unrelated to the original dataset, a goal well aligned with our CSAI classification needs. We do find, however, that because most self-supervised models are trained on object-centric datasets (namely ImageNet [24]), knowledge is not as easily transferable to a scene-centric task and that there are benefits to be had from doing self-supervision with scenes in addition to objects. Our set of contributions is then multi-faceted, from the introduction of indoor scene classification to its application on CSAI data:
@@
    We introduce the novel task of Indoor Scene Classification for CSAI by tailoring data from classic datasets on Scene Classification to yield models useful for CSAI without explicitly training on CSAI data;
@@
    Our thorough evaluation of self-supervised models shows both their usefulness for Indoor Scene Classification but also how further self-supervision on actual indoor scenes improves their downstream performance;
@@
The remaining sections are organized as follows. In Sections II-A and II-B, we review self-supervised and contrastive learning frameworks under deep learning. In Sections II-C and II-D, we bring the most recent works on scene classification and CSAI investigation. In Section III, we describe the datasets, general pipelines used for pretraining, and fine-tuning models used in this work. Then, in Section V, we present the results obtained for our downstream task and real CSAI data. Finally, in Section VI, we point out the contributions, challenges, and possible routes for future work.
@@ II-A Self-supervised Learning
Self-supervised learning (SSL) is a machine-learning approach for building models using unlabeled data. Using unannotated data reduces data annotation costs while finding a purpose to the massive amounts of unlabeled images and texts published daily worldwide [25, 26, 27]. SSL follows a pattern similar to supervised transfer learning with a pre-training step called pretext task, designed to extract features from unlabeled data. In computer vision, common pretext tasks are grayscale colorization [28], jigsaw puzzle, cutout reconstruction, image inpainting, resolution upscaling, foreground object segmentation, clustering, temporal order verification, visual-audio correspondence verification, and contrastive representation comparison.
Afterward, the resulting model is fine-tuned on specific labeled downstream tasks without large amounts of data as a requirement. There are no limitations to the domain of the downstream task. However, it has been empirically shown that ideally the pretext and downstream tasks should comprise images with similar properties, such as textures, shapes, and color gamut [27, 29]. It has been demonstrated that self-supervised models outperform their supervised counterparts for ResNet-50 on image classification, detection, and segmentation in multiple benchmarks [29]. Therefore, SSL shows real advantages over supervised learning, given its higher robustness and generally better performance, especially for CSAI data, which has less than a hundred thousand images in the dataset [30].
@@ II-B Contrastive Learning
Contrastive learning, on the other hand, was thus developed as a new paradigm for self-supervised pre-training [31]. It is closely related to metric learning, an unsupervised learning sub-field that aims to construct a specialized distance metric from weakly supervised data [32].
In the end-to-end contrastive learning framework, two encoders are trained simultaneously with contrastive loss. This method focuses on creating a model that, ideally, compares two samples numerically (e.g., two images), therefore grouping similar images while pushing different ones apart. Within an unlabeled dataset, there is no prior distance measurement. Thus, any augmented version is labeled as “similar” (or positive) for an anchor image sample, while all other samples are labeled as “different” (or negative). The first methods to propose this kind of instance discrimination were IntDisc [33], CMC [34], and Deep InfoMax [35]. However, the approach by [20] was notoriously successful and demonstrated the importance of hard positive samples by introducing a set of 10 data augmentations from which transformations are sampled.
Despite its state-of-the-art performance on several downstream tasks, the end-to-end training is computationally expensive. It does not always lead to the best downstream performance [29]. Thus, many other approaches were developed, such as MoCo [36, 37], SwAV [19], and Barlow Twins [21] to deal with batch size sensitivity; BYOL [38] and SimSiam [39] to overcome the need for negative sampling by changing the instance discrimination pretext task for self-distillation [27]; SupCon [22] to make use of labels in SSL when those are available.
II-C Scene Classification
Humans analyze their environment and important events efficiently, but machines have long stayed caught up in this topic. Scenes are understood as semantically coherent views of real-world environments containing objects, textures, and a spatially separated background [40, 41]. Scene classification is a well-studied field with a particular interest in surveillance, autonomous driving, and robotics navigation, but it is still lacking in performance [42].
Scene classification is closely related to object and texture classification but with multiple objects, textures, backgrounds, and intricate relations between them. Thus, one could imagine it as multiple local classifications happening simultaneously to provide a context and a set of items or actions [43]. However, handcrafting such solution is unfeasible. Then, current approaches [25] use deep learning and scene representation to solve this problem.
@@ II-C Scene Classification
Additionally, [45] used graph nets and image inpainting to show that a scene constitutes minor and essential objects. Only the essential ones matter for the final scene classification, while minor elements can be removed without impact. However, the number of minor objects is usually much higher and can be regarded as noise.
@@
After selecting a pre-trained model, most works develop specialized fine-tuning routines to extract the most scene-centric features in the downstream dataset [43]. GAP-CNN replaces fully connected layers for global average pooling (GAP) to focus on class-specific regions while using fewer parameters [46]; Hierarchical LSTM uses four Long-Short Term Memory (LSTM) modules in an attempt to capture spatial dependencies [47]; DL-CNN proposes dictionary learning (DL) layers to obtain sparse representations and reduce the total number of parameters; finally, the Spatially Unstructured layer was designed to help cope with layout deformations, and scale changes in scene classification [48]. Fisher Vector (FV-CNN) [49] and Mixture of Factor Analyzers Fisher Vector (MFAFVNet) [50] use Improved FV while Multi-scale Orderless Pooling (MOP-CNN) [51] and Semantic Descriptor with Objectness (SDO) [52] employ a Vector of Locally Aggregated Descriptors (VLAD) to cluster local features. However, such an approach heavily increases the number of images to be processed.
@@
Moreover, many studies [53, 54, 55, 56] used object detectors [57, 58, 59, 60, 61] to help with identification of essential scene regions and force the model to pay attention to particular objects during classification. Typically, this dependency introduces two-stage training [62], meaning that learning has to be split between the object detector and classifier.
@@
Several scene classification methods have used multiple layers’ outputs to improve scene representation in this context. The high CNN layers are too compact for the dense level features required for scene classification, keeping only large objects, while the low CNN layers still keep the small ones [63]. Thus, harnessing multi-layer features reaches overall better representations, but careful feature fusion design is required [64]. Ultimately, this is the most structurally complicated approach, and the fused features generate high-dimensional arrays, making the model overfit easily the more layers are used [65].
@@
To the best of our knowledge, few works have used SSL as means to solve scene classification. In scene semantic segmentation, [66] showed an improved model on NYUv2 and SUN RGB-D benchmarks using a model pre-trained on a 5M image 3D rendered dataset called SceneNet RGB-D. In order to tackle the PASCAL VOC benchmark (11,530 images containing 27,450 ROI annotated objects and 6,929 segmentations), Ren and Lee [67] proposed a multi-task learning approach that pre-trains an AlexNet model to predict edges, surface normals, and depth maps from SceneNet RGB-D images and a discriminator to distinguish the synthetic images from the Places365 images. Similarly, She and Xu [68] added a SimCLR contrastive loss as a new pretext task, setting as positive pairs all viewpoints of a single scene. The aforementioned methods reached comparable performance or even outperformed other SSL techniques using the same neural network architecture but underperformed compared to the state-of-the-art method for the tested benchmarks.
@@ II-D CSAI Classification
Easy access and impunity led to a frenzied growth in the consumption and distribution of CSAI over the past few years, with hundreds of millions of images circulating through the web today [2]. The automation of CSAI recognition can protect adults and children, more easily blocking that sort of content in social media and thus preventing trauma [5], but this is still a challenging field of research. Such datasets must stay accessible exclusively to law enforcement personnel, dramatically increasing the difficulties in comparing possible models or performing benchmark evaluations [13]. NCMEC shows that hash comparison tools are already the most significant source for reports today, but law enforcement personnel are the exclusive mechanism responsible for auditing CSAI when dealing with never-before-seen data, which is not only the source of bias but also the psychological strain on the agents’ mental health [8].
@@
In light of these issues, researchers have attempted to design more scalable and reliable methodologies to tackle CSAI recognition. Image hash databases, web crawlers, and file metadata are most commonly used to detect CSAI sources and find criminals [69, 70, 71]. In a more versatile method, NuDetective [72] and iCOP [73] make use of handcrafted nudity image descriptors to build fast CSAI detectors, and [74] used textures and facial distances to distinguish adult from child nudity. However, these solutions are not robust against simple modifications, demanding constant updates to stay functional.
@@
With the recent developments in computer vision detectors, machine learning algorithms also started to be used for CSAI recognition. Ground-up training and transfer learning of CNNs were shown to outperform current forensic and commercial tools easily and can be made into portable tools for search and seizure procedures [9]. Moreover, [10] demonstrated that pornography and age-group recognition could be leveraged for CSAI with substantial gains in accuracy for binary classification, and [8] used a single-model estimation of child presence, age, and gender to improve the performance. [75] arguments apparent age and nudity detection could also be leveraged for CSAI, while [11] focused on non-adult people detection in images for CSAI age estimation, reaching over 90% accuracy on their proposed dataset. When regarding video materials (CSA Videos), studies are fewer and rely on extracting features from individual frames (images) using pre-trained networks and detecting snippets with nudity [73, 76] and extracting facial features to perform victim and abuser identification [77, 78]; Even thought their method was not tested directly on CSAI, [76] used the former approach and took it one step further, incorporating object detection for body parts to rank videos with nudity by “severity”/“harmfulness”.
Figure 1: Methodology pipeline. Self-supervised methods are often used in a common two-stage training protocol: (1) the pretext task — or pre-training stage — and (2) the downstream task — or fine-tuning stage. The pretext task uses unlabeled data and runs the SSL technique, while the downstream task uses labeled data from our downstream task. Our downstream stage is fixed, so we consider the pretext task through the lenses of the three research questions depicted (Section III).
@@
All these studies firmly outline that deep learning can contribute to child abuse inquiries. However, most published works focus on detecting children and nudity/pornography [79, 80, 81] and highlight the difficulties of finding large annotated CSAI databases to experiment with, even with the help of law enforcement agents. Several works focus exclusively on age estimation within CSAI, and children’s images are collected to build datasets that help accomplish such tasks. For instance, [82] proposed VisAGe and [11] AgeMega, databases focused on age estimation with thousands of underage and adult facial images, while [10, 11, 12] gather images from several age estimation databases. Beyond that, [83] and [84] included reproductive organ detection to aid CSAI detection, while [85] used facial expressions to distinguish sensitive from non-sensitive material, which is one of the main challenges in the field. [5] call attention to the fact that “the best results can be achieved if multiple methods are used in combination”, meaning that CSAI is a complex enough problem that age, nudity, and pornography detectors do not fulfill all the roles for possible feature extractors that can aid on this task.
@@
Even with these efforts, CSAI classification presents several challenges before accurate predictions can be made automatically. As [86] explains, the decision of whether an image is explicit enough to be considered indecent is not well defined: the notion of “sexually explicit” may sound straightforward but becomes extremely subtle more often than not, given the context of the image or the culture of the viewers. [87] argue that “nudity alone is not indicative of indecency” whereas posing fully clothed can be considered erotic under some circumstances, mainly if other images depicting the same child in more abusive material are found during the investigation.
@@
Moreover, [13] designed an aggregated analysis pipeline to aid CSAI recognition research by showing a general pipeline for statistical analysis of features, such as scenes, skin tone, and nudity, making use of the region-based annotated child pornography dataset (RCPD) [8], without publicizing information on individual samples. RCPD is a private database used internally by the Brazilian Federal Police and contains 2,138 images among CSAI, pornography, and non-sensitive categories.
@@
In this context, most databases used for CSAI recognition could suffer from biased distributions, not representing what is found in the real world. [13] show that CSA data shared online differs from reported cases of physical abuse in Brazil in terms of race: “most reported victims of child sexual abuse in Brazil are black and brown girls from 8 to 14 years old, but RCPD depicts mostly white children being abused and most of the material apprehended in Brazil has different tendencies than reports of sexual abuse with physical contact”.
@@
Overall, CSAI classification could take advantage of non-direct CSAI recognition, and [2] emphasize that the investigation should optimally consider scene information as means to cluster similar imagery and “report objects that are present, its environment, and potentially identify landmarks that will help locate the region where an abusive image originated”. Experts use scene information to gather context and understand what is happening in the image or video, especially in low-resolution scenarios and when people are covered or with their backs turned [87]. CSAI distribution is a massive problem worldwide, and, for the sake of the children under abuse, combating it demands new data-driven scalable solutions.